40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri May 9 01:36:28 2014, rana, Summary, SUS, Optical Lever QPD Sum trends: they're almost all too weak OLtrend.png
    Reply  Mon May 12 22:42:19 2014, rana, Summary, SUS, Optical Lever QPD Sum trends: they're almost all too weak OL-FB.pngarms_140512.pdf
       Reply  Mon May 12 22:52:59 2014, Jenne, Summary, SUS, Optical Lever Filters are all different OpLevFilters.pdf
       Reply  Tue May 13 13:27:58 2014, Steve, Summary, SUS, Optical Lever Servos setting table OLsums.png
          Reply  Tue May 13 17:03:05 2014, Steve, Summary, SUS, ETMX oplev laser replaced 
             Reply  Tue May 13 17:31:32 2014, Jenne, Summary, SUS, ETMX oplev laser replaced: New oplev gains set OLerr_13May2014.pdfOLfb_13May2014.pdf
                Reply  Tue May 13 22:55:57 2014, Jenne, Summary, SUS, ETMX oplev: cleanup 
             Reply  Wed May 14 10:04:06 2014, Steve, Summary, SUS, oplev laser and temperature TempRules.png
                Reply  Tue May 20 09:12:25 2014, Steve, Summary, SUS, ETMX oplev qpd board ETMXOLQPD121.jpgETMXOLQPDface121.jpgETMXolSUM.png
                   Reply  Tue May 20 15:54:39 2014, rana, Summary, SUS, ETMX oplev qpd board 
Message ID: 9947     Entry time: Tue May 13 17:03:05 2014     In reply to: 9946     Reply to this: 9948   9952
Author: Steve 
Type: Summary 
Category: SUS 
Subject: ETMX oplev laser replaced 

Quote:

Quote:

For some reason or another, I decided that we should see if the optical lever servos were injecting too much noise into the test masses. The ITMs are much worse than the ETMs and I am afeared that they might be making the main noise for our arms in the 20-40 Hz region. Jenne is checking up on these feedback loops to see what's up.

To estimate the actuator gains of the mirrors, I turned on 1 count drives from LSC/CAL oscillators into the LSC drives of each test mass at the frequencies shown in the plot with the resulting peaks showing up in in POX/Y with the single arm locks in red. I will leave these going permanently, but with 0.1 count ampltiudes; we need to make it so in the scripts.

 I'm in the process of filling this table

OPLEV

SERVO     

300 ^

2:0

BR ELP RLP BOOST RES

GAIN

QPD

COUNTS

 

QPD

mW

QPD

beam

OD

HE/NE

output

mW

%

back

on QPD

                                
ETMY PIT  FM1  FM5    55      -30 8,200 0.2   3.3  
        YAW  FM1  FM5    55      -28          
ETMX PIT  FM1  FM5  35        4.4 900 0.2   1.7  
       YAW  FM1  FM5  35       2.1  1,750  0.33    2.8  
ITMY PIT  FM1  FM5        3.3   52 14,400 0.4   9.5  
      YAW  FM1  FM5        3.3  -46          
ITMX PIT  FM1  FM5  50      3.3   30 7,400 0.17   2.8  
       YAW  FM1  FM5  50      3.3  -20          
BS  PIT  FM1  FM5  35      3.3   9 2,800 0.05   2.8  
    YAW  FM1  FM5  35      3.3  -9          
PRM  PIT FM1  FM5  55   FM7  3.3  7 3200 0.06   2.8  
       YAW FM1  FM5  55   FM7  3.3  -5          
SRM  PIT  FM1            -20       9.5  
       YAW  FM1            20          

 I should replace ETMX He/Ne laser

 

ELOG V3.1.3-